# Online Multivariate Regularized Distributional Regression for High-dimensional Probabilistic Electricity Price Forecasting

### v1

@@ 4 Results
Table 5: Computation times. All timings are in seconds. The out-of-sample data for the forecasting study consists of 736 days. We update the 24-dimensional distributional regression model each model on each day. All experiments are run on a standard laptop (Intel Core i7 (16 Threads, 4.9 GHz), 32GB RAM). Estimated speead-ups are calculated by taking Speedup=(Initial $\text{Speedup}=(\text{Initial Fit}\times T)/\text{Total Time}$.
@@ 5 Discussion and Conclusion
Main Results
Our results show that modeling the dependence structure improves forecasting performance and that multivariate distributional regression models yield superior results in terms of the Variogram, Log-Score (LS) and Dawid-Sebastiani Score (DSS) compared to simple benchmark models (LARX). However, we note that multivariate distributional regression models perform better in terms of the LS and DSS than univariate distributional models using a Gaussian copula, while performing worse in terms of the CRPS and Energy Score. This behavior is likely driven by error transmission in the conditional dependence structure, but might also hint at the possibility that the dependence structure in the day-ahead electricity does not change significantly over time. We also find that the Cholesky-based parametrization (which has been used as well in Gioia et al., 2022; Kock and Klein, 2023; Muschinski et al., 2022) of the scale matrix yields better results than the LRA-based parametrization proposed by Salinas et al. (2019). Lastly, we find that the LASSO-regularized estimation of the scale matrix improves forecasting performance significantly. On the computational side, our study can be estimated in roughly 2-3 hours on a standard laptop, providing estimated speed-ups between 80 and 400+ times compared to repeated batch fitting.
Implementation
@@
We implement our algorithm in a fairly generic manner, allowing e.g. for different distributional assumptions and keeping a familiar, sklearn-like API to facilitate the usage by other researchers and data scientists (Pedregosa et al., 2011). We employ just-in-time compilation using numba to further improve the computation speed (Lam et al., 2015). Currently, the code is available on GitHub at https://github.com/simon-hirsch/online-mv-distreg and will be contributed to the ROLCH package (Hirsch et al., 2024).
Future Research
@@
Our research opens multiple avenues for future research. First, further research on the driving forces of the dependence structure in the German electricity market is necessary to improve the forecasting performance and guide decision-making processes in electricity trading. Modeling the dependence structure in electricity markets is a rather open field and has implications beyond forecasting, concerning also risk and portfolio management and asset optimization (Peña et al., 2024; Löhndorf and Wozabal, 2023; Beykirch et al., 2022, 2024). From an algorithmic perspective, we note that while our algorithm is already quite fast, further improvements in the computation speed might be possible by using a CG-type scoring algorithm (Rigby and Stasinopoulos, 2005; Green, 1984; Cole and Green, 1992) and parallelizing over the elements of the distribution parameter. A further open issue is model selection - while the regularized online estimation is fast, the models are still quite complex and can be prone to overfitting. Lastly, due to the generic nature of our implementation, the usage for other high-dimensional forecasting problems such as probabilistic wind, solar and load forecasting can be explored.
@@ Acknowledgments
Simon Hirsch is employed as an industrial PhD student by Statkraft Trading GmbH and gratefully acknowledges the support and funding received. This work contains the author’s opinions and does not necessarily reflect Statkraft’s position. Simon Hirsch is grateful for helpful discussions with Florian Ziel and Daniel Gruhlke.
@@
During the preparation of this work the author used generative AI tools such as ChatGTP and GitHub Copilot in order to improve the quality of the language and of the code. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication
@@ A.1 Abbreviations
A.2 Hyperparameters for the Multivariate Distributional Regression Model
  • Information criteria and model selection: We use the AIC for the multivariate distributional regression model and run the online coordinate descent on an exponential grid of 50 $\lambda$ values. We employ fast model selection based on the first derivatives for the CD-based models.
  • Link functions: We use the identity link for the location for all models. For the CD-based distributional models, we use the InverseSoftPlusLink. For the LRA-based models, we employ the SqrtLink for the diagonal matrix $\bm{\mathrm{A}}$ as initial experiments showed a more robust convergence behavior and the IdentityLink for the matrix $\bm{\mathrm{V}}$. For the degrees of freedom $\nu$, we employ the LogShiftTwoLink, which ensures that $\nu>2$ and hence the covariance matrix is positive definite.
  • Early stopping: We employ early stopping for the path-based regularization of the scale matrix if the AIC does not improve, as described in Section 2.5. We limit the number of off-diagonals for the CD-based parameterization to max 6, however note that the algorithm breaks after fitting 2-3 off-diagonals. We do not limit the number of columns fitted in the LRA-based model and note that the algorithm breaks after fitting the full rank-2 matrix $\bm{\mathrm{V}}$.
A.3 Derivation of Equation 10 and 11 for Newton-Raphson Scoring
@@
A.4 Partial derivatives of the multivariate Gaussian Distribution
@@
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Sigma}})=\frac{1}{(2\pi)^{D/2}}|\bm{\mathrm{\Omega}}|^{1/2}\exp\left(-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Omega}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)$ (36)
@@
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Sigma}})=-\frac{D}{2}\log(2\pi)-\frac{1}{2}\log(|\bm{\mathrm{\Sigma}}|)-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Sigma}}^{-1}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})$ (37)
@@
For the low-rank approximation, we parameterize Equation 36 in terms of the low-rank approximation $\bm{\mathrm{\Omega}}=\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}}$, which yields:
@@
A.5 Partial derivatives of the multivariate t-distribution
@@
and plugging in the according partial derivatives in Equations 54 to 57 and applying integration by parts for the remainder of Equation 53, we have:
@@
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1},\nu)=\log\left(\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}\right)+\log\left(|(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})|\right)+\\ \log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}\right)$ (67)
@@
and hence integration by parts again gives us, similiar to the partial derivatives for the multivariate normal distribution in Equation 48 and 49:

### v2

@@ 4.4 Computation Time
Table 6: Computation times. All timings are in seconds. The out-of-sample data for the forecasting study consists of 1831 days. We update the 24-dimensional distributional regression model each model on each day. All experiments are run on a standard laptop (Intel Core i7 (16 Threads, 4.9 GHz), 32GB RAM). Estimated speed-ups are calculated by taking $\text{Speedup}=(\text{Initial Fit}\times T)/\text{Total Time}$. The figure on the right shows the share of the time spend on the initial fit and on the out-of-sample updating.
Table 6 gives computation times for all experiments. The initial fit for the multivariate distributional regression model takes a few minutes, the update algorithm can be executed in seconds. Using online estimation methods, the experiments can be run in about a few hours on a standard laptop. An estimate for the benefit of online vs. repeated batch fitting can be achieved by multiplying the initial fit duration with 1831 days of out of sample and comparing this to the total time of the online study:
$\text{Speedup}=\frac{\text{Initial Fit}\times T}{\text{Total Time}}.$
By this (albeit simple) measure, the online learning improves computation by a factor of 80 to 600. These estimates are in line with benefits reported in Hirsch et al. (2024) for the univariate online distributional regression case in direct comparison between online estimation and repeated batch estimation.
Figure 16: Efficient frontier of computation time against average log-score for the online multivariate distributional regression model for different update frequencies in the batch and online case. Note that the $x$-axis is in log-scale.
To further analyze the trade-off between computation time and accuracy, we take the OLS-estimated distributional regression model and estimate the model every 1, 7, 14, 30, 60, 180, 365 days online (using a mini-batch online update) and, vice versa, re-estimate the full model in a repeated batch fitting every 7, 14, …, 365 days, using the first subsample of the test data. [^4] Note that already estimating the batch model every 7 days takes more than 10 hours and hence we did not run the experiment for daily estimation in the batch case. We compare the results for online learning, rolling and expanding window batch estimation in Figure 16. For online and batch estimation, we see that increasing update frequency increases forecasting accuracy, but also increases computation times. Crucially, the “efficient frontiers“ of both approaches never intersect. We see that for low computation time budgets, online learning approaches give strictly better results than repeated batch fitting approaches and only for large computation time budgets, expanding window batch estimation takes the lead. A second important observation is that the online learning algorithm is almost constant in the time per update, since we always process a fixed amount of data points. On the other hand the number of data points in the expanding window batch estimation increases, which increases the computation time per fit. Interestingly, the rolling window estimation scheme, which is often favored in the EPF literature, delivers worse predictive accuracy than an expanding window scheme. These results further confirm the estimated speed-up in Table 6.
Summarizing the results on the computational effort, our online update algorithm makes the approach practically viable for researchers and data scientists without access to specialized high-performance computation centers. Taking into account the trade-off between accuracy and computational effort is of practical importance in many industrial settings, where analysts work under time pressure and data arrives at high velocity. Furthermore, time saved in raw computation can be used for better data exploration and feature engineering.
@@ 5 Discussion and Conclusion
Our results show that modeling the dependence structure in the day-ahead market improves probabilistic forecasting performance significantly. First, we see that calibration for prediction bands of the 24-dimensional price path is improved significantly by modeling the dependence structure and moving towards proper joint prediction bands (JPB, see Staszewska-Bystrova, 2011) instead of marginal, quantile-based prediction bands. The online, multivariate distributional regression models deliver strong predictive accuracy across a range of multivariate scoring rules. Additionally, we like to highlight the importance of regularization to avoid overfitting in a high-dimensional setting and conduct two experiments to validate the regularization of the scale matrix. Distributional regression models are interpretable and therefore allow us to discuss the economic interpretation of the time-varying dependence structure.
@@
We analyze the trade-off between computation time and forecasting accuracy. Building an efficient frontier between accuracy and computation time we show that online learning, compared to repeated batch fitting, yields better results for given computation budget — with speed-ups of 2-3 orders of magnitude. Our algorithm is implemented in a fairly generic manner, allowing e.g. for different distributional assumptions and keeping a familiar, sklearn-like API to facilitate the usage by other researchers and data scientists (Pedregosa et al., 2011) and contributed the implementation to the ondil package (Hirsch et al., 2024). [^5] See: https://github.com/simon-hirsch/ondil Reproduction code for all experiments is available on GitHub. [^6] See: https://github.com/simon-hirsch/online-mv-distreg.
@@
Our research opens multiple avenues for future work. First, further research on the driving forces of the dependence structure in the German electricity market is necessary to improve the forecasting performance and guide decision-making processes in electricity trading. Modeling the dependence structure in electricity markets is a rather open field and has implications beyond forecasting, concerning also risk and portfolio management and asset optimization (Peña et al., 2024; Löhndorf and Wozabal, 2023; Beykirch et al., 2022, 2024). From an algorithmic perspective, we note that while our algorithm is already quite fast, further improvements in the computation speed might be possible by using a CG-type scoring algorithm (Rigby and Stasinopoulos, 2005; Green, 1984; Cole and Green, 1992) and parallelizing over the elements of the distribution parameter. A further open issue is model selection - while the regularized online estimation is fast, the models are still quite complex and can be prone to overfitting. Lastly, due to the generic nature of our implementation, the usage for other high-dimensional forecasting problems such as probabilistic wind, solar, and load forecasting can be explored.
@@ Acknowledgments
Simon Hirsch is employed as an industrial PhD student by Statkraft Trading GmbH and gratefully acknowledges the support and funding received. This work contains the author’s opinions and does not necessarily reflect Statkraft’s position. Simon Hirsch is grateful for helpful discussions with Florian Ziel, Daniel Gruhlke, Christoph Hanck and the participants of the IWSM 2025 in Limerick.
@@
Data Statement
The data used in this paper has been provided by Lipiecki et al. (2024) and cannot be shared further by the author. Therefore, for any requests with respect to the data, please contact Lipiecki et al. (2024) directly. A shorter data set of the same explanatory variables is available at the GitHub repository of Marcjasz et al. (2023). [^7] See: https://github.com/gmarcjasz/distributionalnn. For the differences between both data sets please consult Lipiecki et al. (2024), Section 2 and Section 5.1.3.
@@
During the preparation of this work the author used generative AI tools such as ChatGTP and GitHub Copilot in order to improve the quality of the language and of the code (esp. automatic PR reviews, documentation, creation of figures). After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication
@@ A.1 Abbreviations
CP Conformal Prediction
CRPS Continuous Ranked Probability Score
@@
A.2 Derivation of Equation 10 and 11 for Newton-Raphson Scoring
@@
A.3 Partial derivatives of the multivariate Gaussian Distribution
@@
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Omega}})=\frac{1}{(2\pi)^{D/2}}|\bm{\mathrm{\Omega}}|^{1/2}\exp\left(-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Omega}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)$ (35)
@@
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Omega}})=-\frac{D}{2}\log(2\pi)-\frac{1}{2}\log(|\bm{\mathrm{\Omega}}|)-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Omega}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})$ (36)
@@
For the modifed Cholesky-decomposition $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}\bm{\mathrm{D}}^{-1}\bm{\mathrm{L}}$ , the log-likelihood reads:
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{L}},\bm{\mathrm{D}}^{-1})=-\frac{D}{2}\log(2\pi)+\log(|\bm{\mathrm{D}}^{-1}|)-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\left(\bm{\mathrm{L}}^{\top}\bm{\mathrm{D}}^{-1}\bm{\mathrm{L}}\right)(\bm{\mathrm{y}}-\bm{\mathrm{\mu}}).$ (42)
The first derivatives with respect to the elements of $\bm{\mathrm{\mu}}$, $\bm{\mathrm{L}}$ and $\bm{\mathrm{D}}^{-1}$ are given by:
$\frac{\partial\ell}{\partial\bm{\mathrm{L}}_{ij}}$ $=2\bm{\mathrm{D}}^{-1}_{ii}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}$ (44)
$\frac{\partial\ell}{\partial(\bm{\mathrm{D}}^{-1}_{ii})}$ $=\frac{1}{2}\frac{1}{(\bm{\mathrm{D}}^{-1}_{ii})}+\left(\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}\right)^{2}=\frac{1}{2}\bm{\mathrm{D}}_{ii}+\left(\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}\right)^{2}$ (45)
and the second derivatives are given by:
$\frac{\partial\ell^{2}}{\partial\bm{\mathrm{L}}_{ij}^{2}}$ $=-\bm{\mathrm{D}}^{-2}_{ii}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})^{2}$ (47)
$\frac{\partial\ell^{2}}{\partial(\bm{\mathrm{D}}^{-1})_{ii}^{2}}$ $=-\frac{1}{2}\frac{1}{(\bm{\mathrm{D}}^{-1}_{ii})^{2}}=-\frac{1}{2}\bm{\mathrm{D}}_{ii}^{2}$ (48)
keeping in mind that the inverse of a diagonal matrix is given by the inverse of the diagonal elements. For the low-rank approximation, we parameterize Equation 35 in terms of the low-rank approximation $\bm{\mathrm{\Omega}}=\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}}$, which yields:
@@
A.4 Partial derivatives of the multivariate t-distribution
@@
and plugging in the according partial derivatives in Equations 60 to 63 and applying integration by parts for the remainder of Equation 59, we have:
@@
For the MCD, we parameterize $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}$ and hence the PDF is given by:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{L}},\bm{\mathrm{D}}^{-1},\nu)=\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}|\bm{\mathrm{D}}^{-1}|^{1/2}\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}$ (72)
and the log-likelihood is given by
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{L}},\bm{\mathrm{D}}^{-1},\nu)=\log\left(\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}\right)+\log\left(|\bm{\mathrm{D}}^{-1}|\right)+\\ \log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}\right)$ (73)
where $z$ is now defined as
$z=(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}}).$ (78)
For the partial derivatives with respect to the elements of $\bm{\mathrm{D}}^{-1}$, we note that the partial derivatives of the second term are given by:
$\frac{\partial}{\partial\bm{\mathrm{D}}^{-1}_{ii}}$ $=-\frac{1}{2}\bm{\mathrm{D}}_{ii}$ (79)
$\frac{\partial}{\partial(\bm{\mathrm{D}}^{-1}_{ii})^{2}}$ $=\frac{1}{2}\bm{\mathrm{D}}_{ii}^{2}$ (80)
and the partial deriviatives of the third term are given by
$\frac{\partial z}{\partial\bm{\mathrm{D}}^{-1}_{ii}}$ $=\left(\sum_{j=1}^{D}\bm{\mathrm{L}}_{ji}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}$ (81)
$\frac{\partial z}{\partial(\bm{\mathrm{D}}^{-1}_{ii})^{2}}$ $=0$ (82)
$\frac{\partial z}{\partial\bm{\mathrm{L}}_{ij}}$ $=2\bm{\mathrm{D}}^{-1}_{ii}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}$ (83)
$\frac{\partial z}{\partial(\bm{\mathrm{L}}_{ij})^{2}}$ $=2\bm{\mathrm{D}}_{ii}^{-1}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})^{2}$ (84)
Plugging these results in Equations 64 and 65, we have for the partial derivatives of the log-likelihood with respect to the elements of $\bm{\mathrm{D}}$:
$\frac{\partial l}{\partial\bm{\mathrm{D}}^{-1}_{ii}}$ $=-\frac{1}{2}\bm{\mathrm{D}}_{ii}+\frac{(D+\nu)}{2(z+\nu)}\left(\left(\sum_{j=1}^{D}\bm{\mathrm{L}}_{ji}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}\right)$ (85)
$\frac{\partial^{2}l}{\partial(\bm{\mathrm{D}}^{-1}_{ii})^{2}}$ $=\frac{1}{2}\bm{\mathrm{D}}_{ii}^{2}-\frac{(D+\nu)\left(-\left(\left(\sum_{j=1}^{D}\bm{\mathrm{L}}_{ji}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}\right)^{2}\right)}{2(z+\nu)^{2}}$ (86)
and with respect to the elements of $\bm{\mathrm{L}}$:
$\frac{\partial l}{\partial\bm{\mathrm{L}}_{ij}}$ $=\frac{(D+\nu)}{2(z+\nu)}\left(2\bm{\mathrm{D}}^{-1}_{ii}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}\right)$ (87)
$\frac{\partial^{2}l}{\partial(\bm{\mathrm{L}}_{ij})^{2}}$ $=-\frac{(D+\nu)\left((z+\nu)2\bm{\mathrm{D}}_{ii}^{-1}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})^{2}\left(2\bm{\mathrm{D}}^{-1}_{ii}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}\right)^{2}\right)}{2(z+\nu)^{2}}.$ (88)
@@
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{D}},\bm{\mathrm{V}},\nu)=\log\left(\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}\right)+\log\left(|(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})|\right)+\\ \log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}\right)$ (90)
@@
and hence integration by parts again gives us, similiar to the partial derivatives for the multivariate normal distribution in Equation 54 and 55:
@@
A.5 Hyperparameters for the Multivariate Distributional Regression Model
  • 1.
    Information criteria and model selection: We use the BIC for the multivariate distributional regression model and run the online coordinate descent on an exponential grid of 100 $\lambda$ values. We employ fast model selection based on the first derivatives for the CD-based models.
  • 2.
    Link functions: We use the identity link for the location for all models. For the Cholesky-based distributional models, we use the log-link. For the LRA-based models, we employ the square root link for the diagonal matrix $\bm{\mathrm{A}}$ as initial experiments showed a more robust convergence behavior and the identity for the matrix $\bm{\mathrm{V}}$. For the degrees of freedom $\nu$, we employ an inverse softplus shifted to 2.1, which ensures that $\nu>2$ and hence the covariance matrix is positive definite. We have found the shift to be important to avoid numerical instabilities for $\nu$ close to 2. These links also apply to the univariate distributional regression models.
  • 3.
    Early stopping: We employ early stopping for the path-based regularization of the scale matrix if the AIC does not improve, as described in Section 2.6. We limit the number of off-diagonals for the CD-based parameterization to max 6, however note that the algorithm breaks after fitting 1-2 off-diagonals. We do not limit the number of columns fitted in the LRA-based model and note that the algorithm breaks after fitting the full rank-2 matrix $\bm{\mathrm{V}}$.
  • 4.